# On schurity of the dihedral group $\boldsymbol{D_{2p}}$

### v3

@@ 3. Difference sets
For the general theory of difference sets, we refer the readers to [2] and [3, Chapter VI].
@@
Let $p$ be a prime of the form $p=2q+1$, where $q$ is an odd prime, and $D$ a nontrivial difference set in a cyclic group of order $p$. Then $D$ has Paley parameters $(2q+1,q,\frac{q-1}{2})$ or $(2q+1,q+1,\frac{q+1}{2})$.
@@
Eq. (3) implies that $k(k-1)=2q\lambda$. Therefore $k$ or $k-1$ is divided by $q$. Since $D$ is nontrivial, we obtain that $k\in\{q,q+1\}$. Now the required follows from Eq. (3). ∎
Note that $p\equiv 3\mod~{}4$ in Lemma 3.1 because $q$ is odd. For each prime $p$ of the form $p=2q+1$, where $q$ is an odd prime, there is a Paley difference set in a cyclic group of order $p$ with parameters $(2q+1,q,\frac{q-1}{2})$ (see, e.g., [3, Section VI.8]).
Let $G$ be abelian. Following [13], we say that an integer $t$ coprime to $v$ is a (numerical) multiplier of $D$ if $D^{(t)}=\{d^{t}:~{}d\in D\}=Dg=\{dg:~{}d\in D\}$ for some $g\in G$. The set of automorphisms $\sigma_{t}:g\rightarrow g^{t}$, $g\in G$, where $t$ runs over all multipliers of $D$, forms a subgroup of $\operatorname{Aut}(G)$ denoted by $M(D)$. For more information on multipliers of difference sets, we refer the readers to [13].
@@
Let $A$ be a cyclic group of order $p$. Assume that there is a nontrivial difference set $D$ in $A$. Let $(p,k,\lambda)$ be the parameters of $D$. We may assume that $k<\frac{p}{2}$. Indeed, for otherwise $A\setminus D$ is a difference set with $|A\setminus D|<\frac{p}{2}$ and we replace $D$ by $A\setminus D$. If $p$ is a Fermat prime, then $p-1$ is a $2$-power. Since $k$ and $k-1$ are coprime, Eq. (3) yields that $k=p$ or $k=p-1$, a contradiction to nontriviality of $D$.
Now let $p=4q+1$, where $q>3$ is prime.
@@
Eq. (3) implies that $k(k-1)=4q\lambda$. Therefore $k$ or $k-1$ is divided by $q$. Since $D$ is nontrivial and $k<\frac{p}{2}$, we conclude that
@@
Let $M=M(D)$. Clearly, $M\leq\operatorname{Aut}(A)\cong C_{p-1}=C_{4q}$ and hence $|M|$ divides $4q$. Suppose that $|M|$ is even. Then $M$ contains an automorphism of order $2$, i.e. $\sigma_{0}:a\rightarrow a^{-1}\in\operatorname{Aut}(A)$. However, this contradicts to [2, Theorem 3.3] which states that $-1$ is never a multiplier of a nontrivial difference set in a cyclic group. Therefore, $|M|$ is odd and hence
@@
Suppose that $|M|=1$, i.e. $M$ is trivial. Let $t$ be a positive integer such that $t<p$ and the automorphism $\sigma_{t}:a\rightarrow a^{t}$ of $A$ is of order $q$. We are going to prove that $t$ is a nontrivial multiplier of $D$ and by that to obtain a contradiction. Let $n=k-\lambda$. From Lemma 3.2 it follows that $n>1$ and $n<p$. The latter yields that $\operatorname{GCD}(p,n)=1$. Let $u$ be a prime divisor of $n$. There exists an automorphism $\sigma_{u}$ of $A$ such that $\sigma_{u}:a\rightarrow a^{u}$ because $u\leq n<p$. If $|\sigma_{u}|$ is even, then the group $\langle\sigma_{u}\rangle$ contains $\sigma_{0}:a\rightarrow a^{-1}$ and hence $u^{i}\equiv-1\mod~{}p$ for some integer $i$. Therefore $u$ is self-conjugate in the sense of [13, p. 7]. Suppose that $|\sigma_{u}|$ is odd. Then $|\sigma_{u}|=q$ because $|\sigma_{u}|$ divides $|\operatorname{Aut}(A)|=p-1=4q$. Together with $|\sigma_{t}|=q$, this implies that $t\equiv u^{i}\mod~{}p$ for some integer $i$. Finally, observe that
@@
because $\lambda<\frac{k}{2}$. Thus, all conditions of [13, Theorem 3.1] hold for $n_{1}=n$ and $t$ and hence $t$ is a multiplier of $D$.
Now suppose that $|M|=q$. From [3, VI.Theorem 2.6] it follows that there exists a translate of $D$ which is $M$-invariant. Each translate of $D$ is also a difference set in $A$ with the same parameters as $D$ because $A$ is abelian. So without loss of generality, we may assume that $D$ is $M$-invariant. This means that $D$ is a union of some orbits of $M$. Each non-identity orbit of $M$ is a cyclotomic class of order $4$ (see [3, p. 357] for the definition). Recall that $|D|=k\in\{q,q+1\}$ by Lemma 3.2. If $|D|=q$, then $D$ is a non-identity orbit of $M$; if $|D|=q+1$, then $D$ is a union of a non-identity orbit of $M$ with $\{e\}$.
@@
where $x$ and $y$ are integers, by [8, Eq. (51)]. Due to [8, Eq. (55)], we have
@@
If $|D|=q$, then $(1,0)_{4}=\frac{q-1}{4}$ by [3, VI.Theorem 8.10(a)]. Together with Eq. (5), this implies that $x=2q-1-8(1,0)_{4}=2q-1-2q+2=1$. Therefore $p=4q+1=4y^{2}+1$ by Eq. (4) and hence $q=y^{2}$, a contradiction to primality of $q$. If $|D|=q+1$, then $(1,0)_{4}=\frac{q+1}{4}$ by [3, VI.Theorem 8.10(b)]. Together with Eq. (5), this implies that $x=2q-1-8(1,0)_{4}=2q-1-2q-2=-3$. Eq. (4) yields that $p=4q+1=4y^{2}+9$ and hence $q=y^{2}+2$. If $y^{2}\equiv 1\mod~{}3$, then $q$ is divided by $3$. So $q=3$, a contradiction to the assumption of the theorem. If $y^{2}$ is divided by $3$, then $q\equiv 2\mod~{}3$. Therefore $p=4q+1$ is divided by $3$, a contradiction to primality of $p$. ∎
@@
If $q=3$, then there is a difference set in $A$ with parameters $(4q+1,q+1,\frac{q+1}{4})=(13,4,1)$ which is a Singer difference set (see [2]).
Let $G$ be a dihedral group, $A$ a cyclic subgroup of $G$ of index $2$, and $b$ an involution from $G\setminus A$. Suppose that $D$ is a difference set in $A$. Then [26, Lemma 5.1] implies that the partition of $G$ into sets

### v4

@@ 3. Difference sets
For the general theory of difference sets, we refer the readers to [3] and [4, Chapter VI]. The lemma below provides two families of difference sets in cyclic groups. In the present paper, only an existence of such difference sets in cyclic groups is essential for us. A detailed description of the constructions can be found, e.g., in [4, Chapter VI].
@@
Let $A$ be a cyclic group of order $n$.
  1. (1)
    If $n$ is a prime such that $n\equiv 3\bmod 4$, then $A$ has a Paley difference set with parameters $(n,\frac{n-1}{2},\frac{n-3}{4})$.
  2. (2)
    If $n=\frac{q^{d+1}-1}{q-1}$ for some prime power $q$ and positive integer $d\geq 2$, then $A$ has a Singer difference set with parameters $(\frac{q^{d+1}-1}{q-1},\frac{q^{d}-1}{q-1},\frac{q^{d-1}-1}{q-1})$.
Lemma 3.2.
Let $p$ be a prime of the form $p=2q+1$, where $q$ is an odd prime, and $D$ a nontrivial difference set in $C_{p}$. Then $D$ has parameters $(2q+1,q,\frac{q-1}{2})$ or $(2q+1,q+1,\frac{q+1}{2})$.
@@
Eq. (3) implies that $k(k-1)=2q\lambda$. Therefore $k$ or $k-1$ is divisible by $q$. Since $D$ is nontrivial, we obtain that $k\in\{q,q+1\}$. Now the required follows from Eq. (3). ∎
Note that $p\equiv 3\bmod 4$ in Lemma 3.2 because $q$ is odd.
Let $G$ be abelian. Following [17], we say that an integer $t$ coprime to $v$ is a (numerical) multiplier of $D$ if $D^{(t)}=\{d^{t}:~{}d\in D\}=Dg=\{dg:~{}d\in D\}$ for some $g\in G$. The set of automorphisms $\sigma_{t}:g\rightarrow g^{t}$, $g\in G$, where $t$ runs over all multipliers of $D$, forms a subgroup of $\operatorname{Aut}(G)$ denoted by $M(D)$. Further for convenience, we will use the term “multiplier” to call the both integer $t$ and automorphism $\sigma_{t}$. For more information on multipliers of difference sets, we refer the readers to [17].
Lemma 3.3.
[3, Theorem 3.3] The integer $-1$ is never a multiplier of a nontrivial difference set in a cyclic group.
@@
Let $A$ be a cyclic group of order $p$ and $a$ a generator of $A$. Assume that there is a nontrivial difference set $D$ in $A$. Let $(p,k,\lambda)$ be the parameters of $D$. We may assume that $k<\frac{p}{2}$. Indeed, for otherwise $A\setminus D$ is a nontrivial difference set with $|A\setminus D|<\frac{p}{2}$ and we replace $D$ by $A\setminus D$. If $p$ is a Fermat prime, then $p-1$ is a $2$-power. Since $k$ and $k-1$ are coprime, Eq. (3) yields that $k=p$ or $k=p-1$, a contradiction to nontriviality of $D$.
Now let $p=4q+1$, where $q>3$ is a prime.
@@
Eq. (3) implies that $k(k-1)=4q\lambda$. Therefore $k$ or $k-1$ is divisible by $q$. Since $D$ is nontrivial and $k<\frac{p}{2}$, we conclude that
@@
Let $M=M(D)$. Clearly, $M\leq\operatorname{Aut}(A)\cong C_{p-1}=C_{4q}$ and hence $|M|$ divides $4q$. Suppose that $|M|$ is even. Then $M$ contains an automorphism of order $2$, i.e. $\sigma_{0}:a\rightarrow a^{-1}\in\operatorname{Aut}(A)$. However, this contradicts to Lemma 3.3. Therefore, $|M|$ is odd and hence
@@
Suppose that $|M|=1$, i.e. $M$ is trivial. Let $t$ be a positive integer such that $t<p$ and the automorphism $\sigma_{t}:a\rightarrow a^{t}$ of $A$ is of order $q$. We are going to prove that $t$ is a nontrivial multiplier of $D$ and by that to obtain a contradiction. Let $n=k-\lambda$. From Lemma 3.4 it follows that $n>1$ and $n<p$. The latter yields that $\operatorname{GCD}(p,n)=1$. Let $u$ be a prime divisor of $n$. There exists an automorphism $\sigma_{u}$ of $A$ such that $\sigma_{u}:a\rightarrow a^{u}$ because $u\leq n<p$. If $|\sigma_{u}|$ is even, then the group $\langle\sigma_{u}\rangle$ contains $\sigma_{0}:a\rightarrow a^{-1}$ and hence $u^{i}\equiv-1\bmod p$ for some integer $i$. Therefore $u$ is self-conjugate modulo $p$ in the sense of [17, p. 7]. Suppose that $|\sigma_{u}|$ is odd. Then $|\sigma_{u}|=q$ because $|\sigma_{u}|$ divides $|\operatorname{Aut}(A)|=p-1=4q$. Together with $|\sigma_{t}|=q$, this implies that $t\equiv u^{i}\bmod p$ for some integer $i$. Finally, observe that
@@
because $\lambda<\frac{k}{2}$ (Lemma 3.4). Thus, all conditions of [17, Theorem 3.1] hold for $n_{1}=n$ and $t$ and hence $t$ is a multiplier of $D$.
Now suppose that $|M|=q$. From [4, VI.Theorem 2.6] it follows that there exists a translate of $D$ which is $M$-invariant. Each translate of $D$ is also a difference set in $A$ with the same parameters as $D$ because $A$ is abelian. So without loss of generality, we may assume that $D$ is $M$-invariant. This means that $D$ is a union of some orbits of $M$. Each non-identity orbit of $M$ is a cyclotomic class of order $4$ (see [4, p. 357] for the definition). Recall that $|D|=k\in\{q,q+1\}$ by Lemma 3.4. If $|D|=q$, then $D$ is a non-identity orbit of $M$; if $|D|=q+1$, then $D$ is a union of a non-identity orbit of $M$ and $\{e\}$.
@@
where $x$ and $y$ are integers, by [10, Eq. (51)]. Due to [10, Eq. (55)], we have
@@
If $|D|=q$, then $(1,0)_{4}=\frac{q-1}{4}$ by [4, VI.Theorem 8.10(a)]. Together with Eq. (5), this implies that $x=2q-1-8(1,0)_{4}=2q-1-2q+2=1$. Therefore $p=4q+1=4y^{2}+1$ by Eq. (4) and hence $q=y^{2}$, a contradiction to primality of $q$. If $|D|=q+1$, then $(1,0)_{4}=\frac{q+1}{4}$ by [4, VI.Theorem 8.10(b)]. Together with Eq. (5), this implies that $x=2q-1-8(1,0)_{4}=2q-1-2q-2=-3$. Eq. (4) yields that $p=4q+1=4y^{2}+9$ and hence $q=y^{2}+2$. If $y^{2}\equiv 1\bmod 3$, then $q$ is divisible by $3$. So $q=3$, a contradiction to the assumption of the theorem. If $y^{2}$ is divisible by $3$, then $q\equiv 2\bmod 3$. Therefore $p=4q+1$ is divisible by $3$, a contradiction to primality of $p$. ∎
@@
If $q=3$, then there is a difference set in $A$ with parameters $(4q+1,q+1,\frac{q+1}{4})=(13,4,1)$ which is a Singer difference set (see Lemma 3.1(2)).
Let $G$ be a dihedral group, $A$ a cyclic subgroup of $G$ of index $2$, and $b$ an involution from $G\setminus A$. Suppose that $D$ is a difference set in $A$. Then [32, Lemma 5.1] implies that the partition of $G$ into the sets